You roll a number cube. Find p(a or b)
Event a: roll a multiple of 2
Evet b: roll an even number

Answer by stanbon(75887)

You roll a number cube. Find p(a or b)
Event a: roll a multiple of 2
Evet b: roll an even number
---------------------
The 2 events are the same event:
you roll 2,4,or 6 in either case.
P(a or b) = P(a) = 3/6 = 1/2
===================================
Cheers,
Stan H.